OpenStudy (anonymous):

Consider the following functions. f(x) = x2 + 6 g(x) = 2x2 - 5 Find the following: (f - g)(x) =

OpenStudy (anonymous):

f(x) - g(x)

OpenStudy (anonymous):

\[(f-g)(x) = f(x)-g(x)=(x^{2}+6)-(2x^{2}-5) = -x^{2}+1\]

OpenStudy (anonymous):

Thanks joe

OpenStudy (anonymous):

I tried -x^2 + 1 on my hw & it said it was the wrong answer..

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

becasue thats what i thought it was too

OpenStudy (anonymous):

hmmm....maybe try typing it like this?: \[-(x^{2})\] Because there is a difference between: \[-x^{2}\] and this \[(-x)^{2}\] it might have thought you put in the latter.

OpenStudy (anonymous):

no, i tried those too.. hmm..

OpenStudy (anonymous):

The answer i gave is definitely correct, i wager my soul on it. Let me see if i can have someone verify.

OpenStudy (anonymous):

haha ok, thank you.

OpenStudy (anonymous):

Bleh, no one is coming, but thats the answer, im 100% sure of it. For the x^2 term you do (1-2) = -1, for the constant you do OMG THERES MY MISTAKE. its (6-(-5)) = 6+5 = 11 Doing math too fast is bad >.> im sry bout that (pls dont take my soul lol)

OpenStudy (anonymous):

haha so it would be -x^2 + 11 ?

OpenStudy (anonymous):

yes, thats right :)

OpenStudy (anonymous):

yupp (: that worked! thank you! you can keep your soul lol

OpenStudy (anonymous):

lol